A bank contains 30 coins, consisting of nickles, dimes, and quarters. There are twice as many nickles as quarters and the remaining coins are dimes. If the total value of the coins is $3.35, what is the number of each type of coin in the bank?
Answer by stanbon(57290)

A bank contains 30 coins, consisting of nickles, dimes, and quarters. There are twice as many nickles as quarters and the remaining coins are dimes. If the total value of the coins is $3.35, what is the number of each type of coin in the bank?
---------------
Let number of quarters be "x": Value of these is 25x cents
Number of nickels is "2x": Value of these is 5(2x)= 10x cents
Number of dimes is "30-x-2x"=30-3x; Value of these is 10(30-3x)=300-30x cents
----------------
EQUATION:
value + value + value = 335 cents
25x + 10x +300-30x=335
5x=35
x=7 (number of quarters)
2x=14 (number of nickels)
30-3x=9 (number of dimes)
Cheers,
Stan H.